A block weighing 2.3 kg is pushed on a horizontal surface by a force of 80 N directed 30o above the horizontal surface.

the coefficient of kinetic friction between the block and the surface is 0.28,

i. What is the friction force between the block and the surface?         

ii. What is the acceleration of the block?                             


1
Expert's answer
2021-05-12T12:16:10-0400

i.


"F_f=\mu N"


"N=(mg-F\sin30°)=(2.3\cdot9.8-80\cdot \sin30°)=-17.46\ (N)" . So, the body will move along the y-axis and the x-axis. Therefore "F_f=0" .


ii.


"F_y=17.46\ (N)" , "a_y=17.46/2.3=7.59\ (m/s^2)"


"F_x=80\cdot\cos30°=69.28\ (N)" , "a_x=69.28/2.3=30.12\ (m/s^2)"


"a=\sqrt{7.59^2+30.12^2}=31.06\ (m/s^2)"